# envname-P envname#1

### v1

@@ 8 Conclusion
  • Cooldown and QAT Fusion Technique. We introduce a novel approach that performs learning rate decay jointly with quantization-aware training, eliminating redundant full-precision updates and achieving significant compute savings.
@@
While we performed experiments with different datasets and hyperparameters, our work still focuses on a specific LLM architecture, and exact results may differ for different model types. However, we expect the overall observed phenomena to be consistent across different architectures.
@@
We identify several research directions worth exploring. First, the relationship between the optimal QAT fraction and pretraining precision remains unknown. This direction is especially interesting with the emergence of 8-bit floating-point training (Peng et al., 2023) and even 4-bit training (Zhou et al., 2025b; Wang et al., 2025). Second, we are interested in how the observed phenomena are preserved across different training stages. Specifically, how does the optimal QAT fraction change when the full-precision training stage incorporates additional stages such as Supervised Fine-Tuning (SFT) (Lee, 2024), Reinforcement Learning (RL) (Rafailov et al., 2023; Chen et al., 2023), or multimodal training? We speculate on these questions in appendix 23.
@@
We report exact hyperparameters and training approaches used in section 3. Additional experimental information that should facilitate reproduction is summarized in appendix 10, 11, and 21.

### v2

@@ 6 Conclusion
  • Cooldown and QAT Fusion Technique. We introduce a novel approach that performs learning rate decay jointly with quantization-aware training, eliminating redundant full-precision updates and achieving significant compute savings. While being practically useful, this method also suggests that modifications to the universally accepted QAT pipeline can further improve QAT efficiency.
@@
While we checked the generality of our formulation by performing experiments with a different dataset and hyperparameters in appendix J, our work still focuses on a specific LLM architecture, and exact results may differ for different model types. However, we expect the overall observed phenomena to be consistent across different architectures and datasets.
@@
We identify several research directions worth exploring. First, the relationship between the optimal QAT fraction and pretraining precision remains unknown. This direction is especially interesting with the emergence of 8-bit floating-point training (Peng et al., 2023) and even 4-bit training (Zhou et al., 2025b; Wang et al., 2025). Second, we are interested in how the observed phenomena are preserved across different training stages. Specifically, how does the optimal QAT fraction change when the full-precision training stage incorporates additional stages such as Supervised Fine-Tuning (SFT) (Ouyang et al., 2022), Reinforcement Learning (RL) (Rafailov et al., 2023; Schulman et al., 2017), or multimodal training? We speculate on these questions in appendix R.
@@
We report exact hyperparameters and training approaches used in appendix A. Additional experimental information that should facilitate reproduction is summarized in appendix B, C, and N.